# Physics-Informed Wireless Imaging with Implicit Neural Representation in RIS-Aided ISAC System

### v1

@@ III-D Comparison with Traditional Imaging Algorithms
The proposed INR-based imager employs DL-based optimization to extract image information from multipath channel responses and is capable of generating super-resolution images using a small number of CSI measurements. In contrast, model-based methods typically require multipath extraction and retrieve images from a single path, often necessitating a large number of measurements while offering limited resolutions [13, 6, 3, 5]. Moreover, INR supports arbitrary differentiable forward models, whereas FT and CS approaches depend on specific model formulations. Additionally, INR-based imaging is achieved via physics-informed loss functions that integrate domain knowledge into learning [9]. INR does not require constructing large datasets or pre-defining sensing targets, making it superior to traditional DL-based methods in terms of data efficiency and generalization [7, 8].
@@
IV Numerical Results
IV-A Experimental Settings
IV-A1 Simulation Scenario
We consider the simulation setup illustrated in Fig. 1. The center subcarrier frequency is set to 3 GHz. The TX and RX are ULAs with $N_{\text{t}}=N_{\text{r}}=8$ antennas, and their center locations are $10\lambda,10\lambda,0]^{\text{T}}$ and $10\lambda,-10\lambda,0]^{\text{T}}$, respectively. The integrated antenna gain $G_{\text{sen}}=4$. The noise power at each RX antenna is set to $P_{\text{n}}=-110$ dBm. The RIS is deployed in the yOz plane with its center located at $0,0,0]^{\text{T}}$, consisting of $50\times 50$ elements with size $\xi_{\text{s}}=\lambda_{0}/2$, resulting in a UPA of side length 1.53 m. We consider a human imaging scenario, in which the ROI is typically assumed to be 2D in the literature [4]. The ROI center is at $D,0,0]^{\text{T}}$ in the yOz plane with dimensions of $2\text{m}\times 2\text{m}$. The ROI is discretized into $100\times 100$ pixels, with each pixel sized $\lambda/5\times\lambda/5$.
IV-A2 Dataset Generation
We employ a human segmentation dataset from TikTok dance videos on Kaggle [16] to simulate human targets in the ROI. 2615 segmented images are extracted, cropped, resized, and converted into grayscale images of size $100\times 100$. Pixel values are normalized to $0,4\pi A^{2}/\lambda^{2}]$, representing the RCS of a voxel with area $A=\lambda/5\times\lambda/5$ [13]. These normalized pixel values reflect the scattering characteristics of the corresponding voxels. Examples from the synthesized dataset are shown in Fig. 3. Note that the ground-truth images are not required by the proposed INR-based imaging method; they are used only to generate CSI measurements and evaluate imaging accuracy.
IV-A3 Training Details
The NN is trained for 5000 epochs with an initial learning rate of $10^{-3}$. The learning rate is reduced by 50% if the validation accuracy does not improve for 50 consecutive epochs, and training stops if no improvement is observed over 200 epochs. The positional encoding dimension is set to $N_{\text{pe}}=256$, and the $\ell_{2}$-norm of the NN gradients is clipped to be below 2. Training is performed using the Adam optimizer on an Nvidia 4090 GPU with the PyTorch platform. The training grid $\bm{\mathcal{G}}$ is set equal to the output image grid $\bm{\mathcal{G}}^{\circ}$.
IV-A4 Performance Evaluation Metrics
@@
(1) Mean Square Error (MSE): Measures the per-voxel error between the predicted $\hat{\bm{\sigma}}$ and ground truth $\bm{\sigma}$:
@@
where $\mu_{\bm{\sigma}}$ ($\mu_{\hat{\bm{\sigma}}}$) and $\theta_{\bm{\sigma}}^{2}$ ($\theta_{\hat{\bm{\sigma}}}^{2}$) are the mean and variance of $\bm{\sigma}$ ($\hat{\bm{\sigma}}$), respectively, and $\theta_{\bm{\sigma}\hat{\bm{\sigma}}}$ is the covariance between them. Constants $c_{1}$ and $c_{2}$ follow the default settings in the skimage package. SSIM ranges from 0 to 1, with higher values indicating better visual similarity. Unlike MSE and PSNR, SSIM captures structural and perceptual fidelity, reflecting human visual perception.
IV-B Results and Discussions
IV-B1 Imaging Results with Different Systems and Training Settings
TABLE I: Imaging results with different system and training settings (unit for time: second).
@@ 0 RIS phase no RIS DFT random random random random
Imaging result (200 epochs) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Imaging result (5,000 epochs) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
We evaluate the effectiveness of the proposed INR-based imaging algorithm under $K=40$ distinct RIS configurations, with $D=40\lambda$. Table I shows that the proposed method performs well, especially when the sinusoidal activation function and positional encoding are used. After 5000 training epochs, the model produces highly accurate images, with SSIM values approaching 0.99. Even trained for 200 epochs in approximately one second, the INR already generates nearly perfect images. In contrast, using the ReLU activation function results in severely distorted images, and omitting positional encoding also degrades PSNR and SSIM performance. Although omitting the RIS simplifies the forward model and slightly speeds up training, the resulting images are ineffective, indicating the importance of diverse CSI measurements enabled by RIS phase control.
Fig. 4 depicts the training and testing PSNRs for the “no RIS” and “ReLU” configurations. Without RIS, training PSNR increases, but testing PSNR remains below 10 dB, implying overfitting and poor generalization due to insufficient image information. For the ReLU-based NN, neither training nor testing PSNR improves, indicating that ReLU fails to capture the high-frequency features essential for image reconstruction, despite having access to sufficient CSI data.
TABLE II: Imaging results using different algorithms. “GT” denotes ground truth.

### v2

@@ III-D Comparison with Traditional Imaging Algorithms
Figure 5: Imaging mechanisms of model-based, data-driven DL, and INR-based methods.
This subsection compares INR with traditional model-based algorithms and data-driven DL techniques, as illustrated in Fig. 5. Traditional data-driven DL methods require an offline training stage to learn the projection function from CSI measurements to images, demanding a large “CSI-image” dataset [19, 20, 21]. Alternatively, traditional model-based algorithms and the proposed INR-based method only involve online operations, extracting the image information based on the constraints posed by the physical model, rather than from the statistical features of numerous “CSI-image” pairs. Instead of being directly supervised by GT images, INR-based imager transforms the NN outputs to the CSI domain, which is indirectly supervised by CSI measurements, as detailed by (19), (20), and (21). Thus, the INR-based method does not require constructing large “CSI-image” training datasets nor requesting that test target images possess the same statistical distribution with training data. This makes INR possess high data efficiency and generalization.
The data requirement of INR is the same as traditional model-based algorithms. However, the direct model-based image optimization Problem (P1) has been transformed to indirect NN parameter optimization Problem (P2) in INR, since the target image is implicitly represented by the NN and cannot be directly optimized. Nevertheless, the continuous image representation using the NN enables INR to generate arbitrary-resolution images. Additionally, DL-based optimization in INR allows it to support arbitrary differentiable forward models, whereas FT and CS approaches depend on predefined, specific model formulations [9, 17, 16, 8, 10, 11, 18]. The above features make the proposed INR-based imager significantly superior to previous techniques, and the detailed comparison is enumerated in Table I.
Nevertheless, additional issues of INR should be addressed. First, the imaging performance of INR strongly depends on the accuracy of the forward physical model, which guides NN parameter optimization. This issue is also critical for traditional model-based imaging methods [17, 9, 16]. However, INR partially mitigates this problem by leveraging the powerful learning capabilities of DL, as discussed in Sec. IV-A. Second, INR requires online NN training for each imaging target, which introduces a delay in producing the results. By contrast, traditional DL methods allow near real-time inference after offline training. However, our simulations show that INR typically completes training within several seconds, which is considerably faster than the hours-long training time required by conventional DL models [19, 20, 21]. Furthermore, the use of prior information can accelerate the NN training process, as introduced in Sec. IV-B.
IV Enhanced Imager in Practical Environments
This section addresses the challenges of multipath interference and dynamic target imaging to enhance the performance of INR in practical environments.
IV-A Imaging under Multipath Interference
In Sec. III, the term ${\mathbf{H}}_{\text{others},k}$ in (10) is considered as a disturbance to the forward physical model, as shown in (15), due to the unknown characteristics of surrounding environments. However, the performance of INR-based imaging deteriorates when the energy of ${\mathbf{H}}_{\text{others},k}$ becomes large. To enable robust ROI imaging under severe multipath interference, we propose a two-step strategy that involves background calibration and residual interference learning.
IV-A1 Background Calibration
Revisiting the sensing channel model in (10), ${\mathbf{H}}_{\text{others},k}$ includes multipath components scattered by background elements, such as TX-background-RX, TX-background-RIS-RX, and TX-background-ROI-RX. Therefore, the multipaths in ${\mathbf{H}}_{\text{sen},k}$ can be divided into ROI-related and ROI-irrelated components. The latter includes paths scattered only by the RIS and background scatterers. To suppress their interference, ROI-irrelated components can be measured when no targets are present in the ROI and subtracted from the CSI measurements $\widehat{\mathbf{H}}_{\text{sen},k}$ during actual imaging process. This approach, known as background calibration [13], results in a residual channel given by
${\mathbf{H}}_{\text{ima},k}=\underbrace{\mathbf{H}_{\text{roi-bg}}}_{\text{unknown}}+\underbrace{\mathbf{H}_{\text{tx-roi-rx}}+\mathbf{H}_{\text{tx-roi-ris-rx},k}+\mathbf{H}_{\text{tx-ris-roi-rx},k}}_{\text{can be modeled}},$ (24)
where $\mathbf{H}_{\text{roi-bg}}$ denotes multipath scattered by both the ROI and background scatterers. The calibrated CSI measurement is
$\widehat{\mathbf{H}}_{\text{ima},k}={\mathbf{H}}_{\text{ima},k}+\mathbf{N}_{\text{ima}},$ (25)
where $\mathbf{N}_{\text{ima}}$ includes both channel estimation and calibration noise. In practice, calibration noise can be significantly reduced by averaging multiple measurements taken in the absence of targets. The final measurement vector for imaging is then reformulated as $\mathbf{y}=\text{vec}(\{\widehat{\mathbf{H}}_{\text{ima},k}\}_{k=1}^{K})$.
IV-A2 Joint Imaging and Multipath Interference Learning
Although most unknown multipath components are removed via calibration, the forward physical model remains partially known, as described in (24). Following prior work [9, 17, 16], the unknown component $\mathbf{H}_{\text{roi-bg}}$ can be approximated as
$\text{vec}({\mathbf{H}_{\text{roi-bg}}})=\mathbf{A}\bm{\sigma},$ (26)
where $\mathbf{A}\in\mathbb{C}^{N_{\text{t}}N_{\text{r}}\times N_{\text{v}}}$ is an unknown matrix equivalently representing environmental multipath propagations. Therefore, the partially known forward model for the $k$-th RIS configuration becomes
$f^{\prime}_{k}(\bm{\sigma},\bm{\omega}_{k},\mathbf{A})=\mathbf{A}\bm{\sigma}+\text{vec}(\mathbf{H}_{\text{tx-roi-rx}}+\mathbf{H}_{\text{tx-roi-ris-rx},k}\\ +\mathbf{H}_{\text{tx-ris-roi-rx},k}).$ (27)
By stacking $\{f^{\prime}_{k}(\bm{\sigma},\bm{\omega}_{k},\mathbf{A})\}_{k=1}^{K}$, we obtain the complete model $f^{\prime}(\bm{\sigma},\bm{\Omega},\mathbf{A})$. To mitigate the influence of the unknown term $\mathbf{H}_{\text{roi-bg}}$, the matrix $\mathbf{A}$ is treated as a learnable parameter, and (P2) is reformulated as
$\text{(P3)}\ (\bm{\theta}^{\star},\mathbf{A}^{\star})=\mathop{\arg\!\min}\limits_{\bm{\theta},\,\mathbf{A}}\ \mu(f^{\prime}(\mathcal{M}_{\bm{\theta}}(\bm{\mathcal{G}}^{\circ}),\bm{\Omega},\mathbf{A}),\mathbf{y})\\ +\rho(\mathcal{M}_{\bm{\theta}}(\bm{\mathcal{G}}^{\circ})),$
which can be solved using the training framework described in Sec. III-C. Under the supervision of this augmented forward model, both the image representation $\mathcal{M}_{\bm{\theta}}$ and the interference term $\mathbf{A}$ are jointly optimized, thus enabling robust imaging in the presence of multipath interference.
IV-B Successive Imaging for Dynamic Targets
INR is considered an effective imaging technique due to the advantages discussed in Sec. III-D. However, the latency caused by online NN training poses a challenge for dynamic target sensing, where a high image update rate is often required. Although prior work has proposed sophisticated strategies to accelerate training [24, 33], we draw inspiration from medical imaging methods [35] and propose a simple yet effective solution by incorporating prior image information into the training process.
Specifically, we assume that the shape and position of the targets within the ROI are static during one imaging procedure (at a single time instant) but evolve smoothly over time instants [41]. This implies a strong temporal correlation between ROI images at adjacent time steps, allowing historical images to serve as valuable priors. Suppose the image at time t1t-1 is represented by $\mathcal{M}_{\bm{\theta}_{t-1}}$. Then, instead of using only a regularization term $\rho(\cdot)$ as in (P2), we embed prior information from $\mathcal{M}_{\bm{\theta}_{t-1}}$ into the imaging formulation at time $t$, yielding
$\text{(P4)}\quad\bm{\theta}_{t}^{\star}=\mathop{\arg\!\min}\limits_{\bm{\theta}_{t}}\mu(f(\mathcal{M}_{\bm{\theta}_{t}}(\bm{\mathcal{G}}^{\circ}),\bm{\Omega}),\mathbf{y};\mathcal{M}_{\bm{\theta}_{t-1}}),$
where the prior image is used to initialize the parameters of the new network. The optimization is then performed following the same procedure as described in Sec. III-C. This prior-based initialization yields two significant benefits. First, imaging speed is improved, since optimization starts from an already close-to-optimal status. Second, imaging quality is enhanced through knowledge transfer from previous reconstructions. As a result, high-quality and low-latency successive imaging for dynamic targets can be achieved.
V Imaging-Augmented Communication
In addition to supporting vertical industry applications through wireless imaging [9, 10], this section discusses how the high-quality imaging results provided by INR can be leveraged to enhance communication performance.
V-A SE Maximization Problem Formulation
We assume that the positions of all elements in Fig. 2 remain static. Consequently, the communication channel $\mathbf{h}_{\text{com}}$ depends only on $\bm{\omega}$ and is written as $\mathbf{h}_{\text{com}}(\bm{\omega})$. The SE, used as the communication performance metric, is defined as
${\text{SE}}(\bm{\omega})=\log_{2}{\left(1+\frac{P_{\text{t}}\left\|\mathbf{h}_{\text{com}}(\bm{\omega})\right\|^{2}_{2}}{\sigma^{2}_{\text{com}}}\right)},$ (28)
where $\sigma^{2}_{\text{com}}$ denotes the noise variance. The RIS phase configuration $\bm{\omega}$ can therefore be optimized to maximize the SE by exploiting environmental information captured through wireless imaging, formulating the following optimization problem:
(P5) $\max\limits_{\bm{\omega}}$ $\left\|\mathbf{h}_{\text{com}}(\bm{\omega})\right\|^{2}_{2}$
s.t. $|\omega_{n_{\text{s}}}|=1,\ \ \forall n_{\text{s}}=1,\ldots,N_{\text{s}}.$
This problem can be solved by the method proposed in [39], as detailed in Appendix A.
V-B Scenario 1: Communication User Inside the ROI
In this scenario, we consider a user at the initial access stage who has not established a communication link with the BS, and the user location is unknown priorly [37]. The user is supposed a human located within the ROI and holding a communication device, such as a mobile phone, as shown in Fig. 2(a). The corresponding channel is denoted as $\mathbf{h}_{\text{com-in}}$, defined in (3). Accordingly, the RIS phase optimization in (P5) is formulated by substituting
$\mathbf{h}_{\text{com}}(\bm{\omega})=\Big(g_{\text{com}}g_{\text{ris}}\mathbf{H}_{\text{tx-ris}}\text{diag}(\mathbf{h}_{\text{ris-ue}})\Big)\bm{\omega}+\mathbf{h}_{\text{tx-ue}}.$ (29)
To optimize RIS phases for high communication performance, the user location should be obtained. However, conventional localization methods depend on user cooperation [44] and may suffer from low estimation accuracy. Moreover, prior studies often model the user as a point target, while in practice, the human body position may differ from the communication device position [38, 37, 50], potentially degrading the beamforming performance, especially in near-field regions.
To address this, the proposed ISAC system first treats the user as a passive target, performing active sensing to reconstruct an image of the ROI using the proposed INR-based methods. Hence, the user simultaneously serves as a communication terminal and a target being sensed. The ROI image is then processed by a hand detection NN to localize the user’s hand position [38, 51]. Consequently, this detected position is assumed to be the communication device location and is subsequently used for RIS phase optimization, thereby ensuring accurate RIS-aided beamformers and reducing beam scanning time [37].
V-C Scenario 2: Communication User Outside the ROI
In this scenario, we consider a user who has established a communication link with the BS, and the user location can be estimated using RIS-aided localization algorithms [44, 52]. The user is assumed to be located outside the ROI, and the communication channel is given by $\mathbf{h}_{\text{com-out}}$, as shown in Fig. 2(b). The ROI involves other targets, which introduce additional multipath components to $\mathbf{h}_{\text{com-out}}$, including $\mathbf{h}_{\text{tx-roi-ue}}$, $\mathbf{h}_{\text{tx-roi-ris-ue}}$, and $\mathbf{h}_{\text{tx-ris-roi-ue}}$, as defined in (6). Although the RIS phases can be optimized for communication without target imaging, the ROI-related multipath components are unknown and cannot be properly considered during RIS phase optimization, potentially reducing the effective energy of $\mathbf{h}_{\text{com-out}}$.
To address this, we propose to employ the INR-based imaging technique to generate a high-quality ROI image, allowing accurate characterization of the ROI-related multipath components. Based on the channel models in (7) and (8), (P5) can be formulated by substituting [23]
$\mathbf{h}_{\text{com}}(\bm{\omega})=g_{\text{com}}g_{\text{ris}}\Big(\mathbf{H}_{\text{tx-ris}}\text{diag}(\mathbf{h}_{1})+\mathbf{H}_{1}\text{diag}(\mathbf{h}_{\text{ris-ue}})\Big)\bm{\omega}\\ +\mathbf{h}_{\text{tx-ue}}+\mathbf{h}_{\text{tx-roi-ue}},$ (30)
where $\mathbf{h}_{1}=\mathbf{h}_{\text{ris-ue}}+\mathbf{H}_{\text{ris-roi}}\text{diag}(\bm{\sigma})\mathbf{h}_{\text{roi-ue}}$, and $\mathbf{H}_{1}=\mathbf{H}_{\text{tx-roi}}\text{diag}(\bm{\sigma})\mathbf{H}_{\text{roi-ris}}$. In this manner, the RIS phase design incorporates both ROI- and RIS-related multipaths, leading to improved SE performance.
VI Numerical Results
VI-A Experimental Settings
VI-A1 Simulation Scenario
We consider the simulation setup illustrated in Fig. 2. The center subcarrier frequency is set to 3 GHz. The TX and RX are ULAs with $N_{\text{t}}=N_{\text{r}}=8$ antennas, and their center locations are $10\lambda,10\lambda,0]^{\text{T}}$ and $10\lambda,-10\lambda,0]^{\text{T}}$, respectively. The integrated antenna gains, i.e., $G_{\text{com}}$ and $G_{\text{sen}}$, are set to 4 [53]. The received noise power at each antenna of the RX is set to $P_{\text{n}}=-110$ dBm [31]. The RIS is deployed in the yOz plane with its center located at $0,0,0]^{\text{T}}$, consisting of $50\times 50$ elements with size $\xi_{\text{s}}=\lambda/2$, resulting in a uniform planar array of side length 2.5 m. We consider a human imaging scenario, in which the ROI is typically assumed to be 2D in literature [14, 9, 38], whereas 3D imaging results are presented in Sec. VI-B7. The center of the ROI is located at $D,0,0]^{\text{T}}$ with dimensions of $2\text{m}\times 2\text{m}$. The ROI lies in the yOz plane and is discretized into $100\times 100$ pixels, with each pixel sized $\lambda/5\times\lambda/5$, which is smaller than the diffraction resolution limit presented in [18].
VI-A2 Data Generation
We employ a human segmentation dataset derived from TikTok dance videos on Kaggle [54] to simulate human targets in the ROI. A total of 2,615 segmented images are extracted, cropped, resized, and converted into grayscale images of size $100\times 100$. Pixel values are normalized to the range $0,4\pi A^{2}/\lambda^{2}]$, representing the radar cross section (RCS) of a voxel with area $A=\lambda/5\times\lambda/5$ [17]. These normalized pixel values reflect the scattering characteristics of the corresponding voxels. Examples from the synthesized dataset are shown in Fig. 6. Note that the GT images are not required by the proposed INR-based imaging method; they are used only to generate CSI and evaluate imaging accuracy in the experiments. The CSI is generated using the ray tracing channel model given in [43], as presented in Sec. II.
@@ VI-A2 Data Generation
VI-A3 Training Details
The NN is trained for 5000 epochs with an initial learning rate of $10^{-3}$. The learning rate is reduced by 50% if the validation accuracy does not improve for 50 consecutive epochs, and training stops if no improvement is observed over 200 epochs. The positional encoding dimension is set to $N_{\text{pe}}=256$, and the $\ell_{2}$-norm of the NN gradients is clipped to be below 2. Training is performed using the Adam optimizer on an Nvidia 4090 GPU with the PyTorch platform. Unless otherwise specified, the training grid $\bm{\mathcal{G}}$ is set equal to the output image grid $\bm{\mathcal{G}}^{\circ}$.
VI-A4 Performance Evaluation Metrics
@@
(1) Mean Square Error (MSE): Measures the per-voxel error between the predicted $\hat{\bm{\sigma}}$ and GT $\bm{\sigma}$:
@@
TABLE II: Imaging results with different system and training settings (unit for time: second).
@@ 0 RIS phase no RIS DFT random random random random
Imaging result (200 epochs) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Imaging result (5,000 epochs) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
where $\mu_{\bm{\sigma}}$ ($\mu_{\hat{\bm{\sigma}}}$) and $\theta_{\bm{\sigma}}^{2}$ ($\theta_{\hat{\bm{\sigma}}}^{2}$) are the mean and variance of $\bm{\sigma}$ ($\hat{\bm{\sigma}}$), respectively, and $\theta_{\bm{\sigma}\hat{\bm{\sigma}}}$ is the covariance between them. Constants $c_{1}$ and $c_{2}$ follow the default settings in the skimage python package. SSIM ranges from 0 to 1, with higher values indicating better visual similarity. Unlike MSE and PSNR, SSIM captures structural and perceptual fidelity, reflecting human visual perception.
VI-B Results and Discussions
VI-B1 Imaging Results with Different Systems and Training Settings
@@
We evaluate the effectiveness of the proposed INR-based imaging algorithm under $K=40$ distinct RIS configurations, with the ROI positioned at $D=40\lambda$. Table II shows that the proposed method performs well, especially when the sinusoidal activation function and positional encoding are used. After 5000 training epochs, the model produces highly accurate images, with SSIM values approaching 0.99. Even at 200 epochs, the INR already generates nearly perfect images, with the cost of time in approximately one second. In contrast, using the ReLU activation function results in severely distorted images, and omitting positional encoding also degrades PSNR and SSIM performance. Random RIS phase configurations outperform the DFT codebook in sensing performance. Although omitting the RIS simplifies the forward model and slightly speeds up training, the resulting images are ineffective, indicating the importance of diverse CSI measurements enabled by RIS phase control.
Fig. 7 illustrates the training and testing PSNRs for the “no RIS” and “ReLU” configurations. Without RIS, training PSNR increases, but testing PSNR remains below 10 dB, implying overfitting and poor generalization due to insufficient image information. For the ReLU-based NN, neither training nor testing PSNR improves, indicating that ReLU fails to capture the high-frequency features essential for image reconstruction, despite having access to sufficient CSI data.
VI-B2 Hyperparameter Discussions
TABLE III: Imaging performance versus $\alpha$.
0 $\alpha$ $10^{0}$ $^{-0.5} 101.0$$^{-1.0}$ $10^{-1.5}$ $10^{-2.0}$
PSNR (dB) 16.6162 21.2241 24.8158 27.1059 29.2841
SSIM 0.7514 0.8756 0.9513 0.9807 0.9889
Figure 8: Imaging performance with varying $\chi$.
This subsection discusses the choice of the hyperparameter $\alpha$ in the loss function and $\chi$ in positional encoding.
First, we have employed a sparsity penalty $\alpha$ in the loss function to stabilize the NN training process and prevent meaningless results. To address the influences of $\alpha$, additional simulations are conducted, and the results are listed in Table III. When $\alpha$ is large, the sparsity penalty term dominates in the loss function, leading to low supervision ability of the physical model. Thus, the imaging performance is degraded. According to Table II, $\alpha=0$ and $\alpha=0.01$ derive approximately the same imaging performance. Consequently, $\alpha=0.01$ is a desirable choice for NN training and has been utilized for subsequent simulations.
Second, the variance of the elements in matrix $\mathbf{B}$ is determined by $\chi^{2}$, influencing Fourier frequencies in positional encoding. According to [48], $\chi$ is chosen for each task with a hyperparameter sweep, and our simulation results with respect to $\chi$ are shown in Fig. 8. In our considered scenarios, $\chi=4$ realizes the best performance of PSNR and SSIM. Thus, this value is employed for the following subsections.
VI-B3 Imaging Result Comparison of INR with FT and CS Algorithms
TABLE IV: Imaging results using different algorithms.